# AiGet: Transforming Everyday Moments into Hidden Knowledge Discovery with AI Assistance on Smart Glasses

### v1

@@ 5. AiGet System
We introduce AiGet, a proof-of-concept system that aligns with the aforementioned design goals (Sec 4.5.3). In this section, we detail the primary features of the AiGet system, its implementation, and an in-lab simulated ablation study conducted to evaluate its knowledge generation ability.
@@ 5.1. Key Features of AiGet System
Figure 3. AiGet’s System Processing Pipeline. Some details are omitted in the figure, and full Agents’ prompts can be found in Appendix C.
@@ 5.1.2. Context Analysis (D1, D2)
To enhance analysis and support D2, the system provides the LLM with a user profile (Xu et al., 2024; Wang et al., 2024b) containing values, interests, and basic demographic information. A sample profile is as follows: “”Values/Interest”: [”healthy life”, ”fitness”, ”flowers”, ”coffee lover”, ”cat”, ”fun facts/history”, ”design”] [^6] To ensure broad coverage of user interests and values, we followed a two-step process: (1) Curating 14 informal knowledge topics (e.g., Animals & Plants) and 58 value categories (e.g., Physical Value-Beauty) from taxonomies in prior literature (Livingstone, 2001a; Marino et al., 2019; Schwartz, 1992; Human Science, [n. d.]) to guide participant selection; (2) Encouraging participants to expand the list with detailed topics (e.g., ”Plant Care Tips”) that aligned with their experiences and learning desires., ”Age”: ”30”, ”Gender”: ”female”, ”Citizenship”: ”XX”, ”Residence”: ”XX”, ”Education”: ”PhD in visual design”, ”Occupation”: ”senior student in XX””. Using this profile, the LLM assesses the user’s familiarity with the current location and surrounding entities (e.g., ”a senior student is familiar with the campus,” or ”coffee lovers know the basics of different types of coffee”). It also refines learning desires and identifies potential connections between the environment and the user’s interests (e.g., linking the topic of ”fruit nutrition” to a user’s interest in a ”healthy lifestyle” when passing the supermarket fruit aisle).
@@
This table summarizes challenges in multimodal entity identification and solutions using MLLM prompt refinements. For Ambiguity in Gaze Interpretation, Context-Guided Selections prioritize rare elements to enhance relevance (e.g., focusing on a museum specimen rather than its jar). For Visual Recognition Errors, Multimodal-Facilitated Visual Processing uses text cues to identify underrepresented entities, such as niche products. These methods improved contextual accuracy and robustness in pilot tests. Challenges Solutions & Steps Ambiguity in Gaze Interpretation: When gaze direction overlaps multiple entities.
@@ 5.1.3. Knowledge Generation and Prioritization (D1, D2, D3)
This table outlines prompt refinements for LLMs to enhance novelty and unexpectedness in content generation. To improve novelty, challenges like generic outputs and repetition were addressed using few-shot prompting tailored to user expertise and a filtering mechanism inspired by retrieval-augmented generation, ensuring the generation of fresh, relevant content. For unexpectedness, prompts incorporated a ”surprise taxonomy,” guiding models to focus on lesser-known facts, misconceptions, and unseen but intriguing perspectives, enhancing engagement and user interest. Aspect Challenges Solutions & Steps Novelty Balancing Detail with Conciseness: Requirement for Conciseness led to generic outputs from naive LLMs, lacking depth for educated users. Example: ”Palm trees play an important role in the ecosystem by providing food for various species.” Few Shot Prompting: Employed few-shot prompting (Reynolds and McDonell, 2021) with concrete examples to guide LLMs in tailoring details to users’ education levels. Example: ”Palm trees emit volatile organic compounds attracting beneficial insects, reducing pesticide use.” balances conciseness with novel details, and domain experts could receive more technical precision. Reducing Repetition when Revisiting Entities: Repetition undermined novelty, especially when revisiting recurring entities like trees along a route. Simply feeding full histories and requiring new content failed due to LLMs’ limitations in long context processing. Filtering Mechanism with Chain of Thoughts: Inspired by retrieval-augmented generation (RAG) (Zamani et al., 2022), only fed the 10 most relevant historical items based on semantic similarity [^8] Using the all-MiniLM-L6-v2 model, https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2 to contextual description. Applied Chain-of-Thought method (Wei et al., 2022) to break complex tasks into smaller steps and maximize the likelihood of generating novel content: Identify entities/topics already mentioned in concise history. Determine unexplored gaps. Generate content focusing on new perspectives or entities. Post-generation, filter content highly similar to history ($\geq 0.75$) to ensure novelty. Unexpectedness Lack of Surprising Perspectives: Hard to generate unexpected content without explicit guidance. Specifying Potential Unexpectedness Category: Guided by ”surprise taxonomy” (Modirshanechi et al., 2022), prompts suggested: Unseen but interesting entities (Observation-Mismatch Surprise (Modirshanechi et al., 2022)). Lesser-Known Fun facts (e.g., trivia) or misconceptions (Belief-Mismatch Surprise (Modirshanechi et al., 2022)).
@@ 5.3. In-Lab Evaluation of AiGet’s Knowledge Generation Pipeline
Specifically, we conducted an ablation study by removing two key components of AiGet pipeline—knowledge generation rules (D3) and personalization (D2)—from the pipeline, following the method used in prior research (Xu et al., 2024; Wang et al., 2024b). The pipeline originally consisted of a Multimodal LLM, User Profile, and Rules (Prioritization with Focus/Familiarity Analysis). We kept the Multimodal LLM (D1) as it is compulsory for context-relevant knowledge generation. Then, we compared the knowledge generated by the resulting three pipelines to understand what components and associated guidelines affect the desirability of knowledge. Note: D4 is not considered here, as it focuses on interaction rather than knowledge generation.
@@ 5.3.3. Study Design and Data Preparation
A within-subject design was used to compare the knowledge generation quality across different pipelines. Each participant rated knowledge generated from 9 moments, with 3 knowledge items per moment (one from each pipeline), totaling 27 knowledge per participant.
@@ 5.3.5. Results
This underscores the need for AI systems to dynamically infer user interests and intentions from context, as users may not explicitly define all their fine-grained desires in advance. It also highlights the importance of coordinating rules with user profiles.
@@ 6.1. Participants
The study consisted of two phases. In Phase 1, 18 participants (8 female, 10 male, Age: $M=23.9,\>SD=3.7$) from the university community engaged in a one-time use of AiGet. Among these participants, six reported using AI tools (e.g., Google Lens, ChatGPT app) to identify surrounding objects in their daily activities at least three times per week. Eleven participants used such tools less frequently, ranging from once per month to twice per week. The remaining participant rarely used these tools. All participants self-reported a minimum level of professional working fluency in English.
In Phase 2, 6 participants (3 female, 3 male) from Phase 1 were selected for continued usage based on their demographic diversity and availability (see Appendix D.2 for details). These participants, representing a range of self-reported engagement with their surroundings—3 with high engagement and 3 with lower engagement—took part in additional experiments involving multiple sessions.
@@ 7. Overall Discussion
In the following discussion, we will highlight two key aspects: 1) how wearable knowledge discovery assistants (i.e., wearable AI assistants supporting knowledge discovery) can be used for informal learning in daily life, and 2) what improvements are needed for long-term usage of such assistants. By examining these aspects, we aim to provide insights into the future development and integration of context-aware wearable knowledge discovery assistants in everyday contexts, further improving how we acquire and interact with information from our in-situ environment.
@@ 7.1.3. Improving User Trust for “Surprising” Content
Enhancing user trust in wearable knowledge discovery assistants is crucial, particularly when the system presents surprising or counter-intuitive information. As P1 noted, prior negative experiences with AI-generated content (e.g., hallucinations (Bang et al., 2023)) can cause users to distrust unexpected information, even when it is accurate—such as learning that ‘a palm tree is more closely related to grass than typical trees’.
@@ 7.2.1. Balancing Knowledge Diversity and Depth
In the real-world study, participants highlighted AiGet’s ability to “break out of information cocoons [exposed only to information that aligns with their existing beliefs or interests] (P3)” by providing information on both prior interested topics and relatively unfamiliar content embedded in the environment, thus expanding their “knowledge base (P18)”. While this feature fosters the discovery of new interests, it also raised concerns about potential information overload. As P1 noted, “I don’t need to be an expert on everything,” emphasizing the need for careful design in content delivery, especially for repeated exposure to entities of limited interest.
@@ 7.2.3. Considering Interruptibility in Real-Time Mode and Supporting Retrospective Mode
To reduce the interference to primary tasks, beyond providing easy-to-digest content and enabling manual output adjustment, the future system could further predict user interruptibility (Pejovic and Musolesi, 2014) and automatically reduce its frequency or turn off during moments when they preferred fewer interruptions (as shown in Figure 2-Gray parts). This feature can be implemented by 1) Using MLLMs for FPV analysis and withholding notifications when users are engaged in high-concentration or time-sensitive activities (e.g., crossing a road, focusing on work, and rushing in train stations); 2) Incorporating additional input modalities, such as Electrodermal Activity signals (Tan et al., 2021), to discern variations in mental or emotional states even within the same activity, allowing nuanced adjustments (e.g., distinguishing whether a student is thinking of homework after class or enjoying a leisurely walk home); and 3) Adapting interruptibility rules based on individual preferences and interaction history. Complementing this real-time adaptation, participants (e.g., P3) suggested incorporating a retrospective mode to catch up on interesting or useful information they might have missed while the system is in “Do Not Disturb” status. This retrospective mode could also encourage reflection on daily experiences, offering meta-cognitive insights excluded in real-time mode due to the high cognitive load involved. Additionally, it could converse with users during this period and dynamically update user preferences in the profile based on their reflections on interactions throughout the day.
@@ 8. Limitations
While our evaluation provided initial insights into using a proactive wearable knowledge discovery assistant, it had several limitations that future research should address. First, the study mainly involved university students and staff, who may represent early adopters of smart glasses technology. Future studies should include a more diverse demographic to ensure broader applicability. Furthermore, although we conducted multi-session studies over a week, longer-term evaluations are necessary to better understand the system’s impact on learning behaviors and retention. Additionally, due to hardware constraints, participants could not use the system continuously in their daily lives without being observed, potentially affecting the naturalness of their interactions. Participants may also have been more receptive to receiving knowledge during study sessions, where they faced fewer cognitive pressures from primary tasks (e.g., casual walking and shopping). Future research should aim for extended, real-world deployments to capture more authentic usage patterns, including user preferences when engaging with the system under varying temporal and cognitive demands, and assess the system’s long-term effects on learning and behavior.
@@ 9. Conclusion
We explored the feasibility of a wearable AI assistant, AiGet, that proactively analyzes users’ in-situ environments to provide personalized, context-relevant, yet often overlooked knowledge with minimal disruption to primary tasks. Through in-lab and real-world evaluations, we demonstrated AiGet’s utility in uncovering overlooked interests, enhancing task enjoyment, reviving curiosity, and deepening connections with the environment. Our findings highlight the potential of AI-assisted informal learning to transform everyday moments into enriching learning experiences. We also identified key design considerations for future systems, including balancing knowledge diversity and depth, enhancing coherence between learning moments, and supporting real-time and retrospective interaction modes. As wearable AI assistants continue to evolve, AiGet represents a promising initial step toward seamlessly integrating informal learning into daily life, opening new avenues for ubiquitous knowledge discovery and personal growth.
@@ A.2. Knowledge Categories and Examples
This table defines four knowledge types with examples from a supermarket beverage scene. Factual Knowledge includes specific details, like product ingredients (e.g., ”Apple juice contains natural antioxidants”). Conceptual Knowledge explains relationships, such as nutritional comparisons (e.g., ”Apple juice is sweeter, orange juice has more vitamin C”). Procedural Knowledge covers methods, like creating a mocktail (e.g., ”Mix apple juice, orange juice, and ginger ale”). Metacognitive Knowledge involves self-reflection, such as understanding personal decision-making criteria. Knowledge Type Definition Examples Factual Knowledge Knowledge of terminology and specific details and elements “This apple juice contains natural antioxidants called polyphenols, which can help protect your cells from damage and may even support heart health”
@@ B.1. FPV Difference Calculation
B.2. Parallel Processing during Output Transformation

### v2

@@ 5. AiGet System
We introduce AiGet, a proof-of-concept system that aligns with the aforementioned design goals (Sec 4.5.3). In this section, we detail the primary features of the AiGet system, its implementation, and an in-lab simulated ablation study conducted to evaluate its knowledge generation capability.
@@ 5.1. Key Features of AiGet System
Figure 3. AiGet’s System Processing Pipeline. (1) LLM Request Trigger: Supports mixed-initiative knowledge queries. (2) Context Analysis: Analyzes real-time user attention and long-term interests and infers learning desires. (3) Knowledge Generation & Prioritization: Filters redundant content, prioritizing the most valuable knowledge to avoid overload. (4) Output Transformation: Presents knowledge in a multimodal format to balance engagement and cognitive load. (5) Follow-Up User Actions: Enables user control over interaction. Some details are omitted in the figure. Full details and agent prompts are in Appendix C.
@@ 5.1.2. Context Analysis (D1, D2)
To enhance analysis and support D2, the system provides the LLM with a user profile (Figure 3 (User Profile)) (Xu et al., 2024; Wang et al., 2024b) containing values, interests [^6] To ensure broad coverage of user interests and values, we followed a two-step process: (1) Curating 14 informal knowledge topics (e.g., Animals & Plants) and 58 value categories (e.g., Physical Value-Beauty) from taxonomies in prior literature (Livingstone, 2001a; Marino et al., 2019; Schwartz, 1992; Human Science, [n. d.]) to guide participant selection; (2) Encouraging participants to expand the list with detailed topics (e.g., ”Plant Care Tips”) that aligned with their experiences and learning desires., and basic demographic information (i.e., age, gender, nationality, residence, and academic/professional background; see Appendix B.2 for a sample profile). Using this profile, the LLM assesses the user’s familiarity with the current location and surrounding entities (e.g., ”a senior student is familiar with the campus,” or ”coffee lovers know the basics of different types of coffee”). It also refines learning desires and identifies potential connections between the environment and the user’s interests (e.g., linking the topic of ”fruit nutrition” to a user’s interest in a ”healthy lifestyle” when passing the supermarket fruit aisle).
@@
This table summarizes challenges in multimodal entity identification and solutions using MLLM prompt refinements. For Ambiguity in Gaze Interpretation, Context-Guided Selections prioritize rare elements to enhance relevance (e.g., focusing on a museum specimen rather than its jar). For Visual Recognition Errors, Multimodal-Facilitated Visual Processing uses text cues to identify underrepresented entities, such as niche products. These methods improved contextual accuracy and robustness in pilot tests.
Challenges Solutions & Steps
Ambiguity in Gaze Interpretation: When gaze direction overlaps multiple entities.
@@ 5.1.3. Knowledge Generation and Prioritization (D1, D2, D3)
This table outlines prompt refinements for LLMs to enhance novelty and unexpectedness in content generation. To improve novelty, challenges like generic outputs and repetition were addressed using few-shot prompting tailored to user expertise and a filtering mechanism inspired by retrieval-augmented generation, ensuring the generation of fresh, relevant content. For unexpectedness, prompts incorporated a ”surprise taxonomy,” guiding models to focus on lesser-known facts, misconceptions, and unseen but intriguing perspectives, enhancing engagement and user interest. Aspect Challenges Solutions & Steps Novelty Balancing Detail with Conciseness: Requirement for Conciseness led to generic outputs from naive LLMs, lacking depth for educated users. Example: ”Palm trees play an important role in the ecosystem by providing food for various species.” Few Shot Prompting: Employed few-shot prompting (Reynolds and McDonell, 2021) with concrete examples to guide LLMs in tailoring details to users’ education levels. Example: ”Palm trees emit volatile organic compounds attracting beneficial insects, reducing pesticide use.” balances conciseness with novel details, and domain experts could receive more technical precision. Reducing Repetition when Revisiting Entities: Content repetition undermined novelty, especially when revisiting recurring entities like trees along a route. Simply feeding full histories and requiring new content failed due to LLMs’ limitations in long context processing. Filtering Mechanism with Chain of Thoughts: Inspired by retrieval-augmented generation (RAG) (Zamani et al., 2022), only fed the 10 most relevant historical items based on semantic similarity* to contextual description. Applied Chain-of-Thought method (Wei et al., 2022) to break complex tasks into smaller steps and maximize the likelihood of generating novel content: Identify entities/topics already mentioned in concise history. Determine unexplored gaps. Generate content focusing on new perspectives or entities. Post-generation, filter content highly similar to history ($\geq 0.75$)* to ensure novelty. Unexpectedness Lack of Surprising Perspectives: Hard to generate unexpected content without explicit guidance. Specifying Potential Unexpectedness Category: Guided by ”surprise taxonomy” (Modirshanechi et al., 2022), prompts suggested: Unseen but interesting entities (Observation-Mismatch Surprise (Modirshanechi et al., 2022)). Lesser-Known Fun facts (e.g., trivia) or misconceptions (Belief-Mismatch Surprise (Modirshanechi et al., 2022)). * The similarity is calculated using the all-MiniLM-L6-v2 model, https://huggingface.co/sentence-transformers/all-MiniLM-L6-v2
@@ 5.3. In-Lab Evaluation of AiGet’s Knowledge Generation Pipeline
Specifically, we conducted an ablation study by removing two key components of AiGet pipeline—knowledge generation rules (D3) and personalization (D2)—from the pipeline, following the method used in prior research (Xu et al., 2024; Wang et al., 2024b). The pipeline originally consisted of a Multimodal LLM, User Profile, and Rules (Prioritization with Focus/Familiarity Analysis). We kept the Multimodal LLM (D1) as it is compulsory for context-relevant knowledge generation. Then, we compared the knowledge generated by the three resulting pipelines to understand what components and associated guidelines affect the desirability of knowledge. Note: D4 is not considered here, as it focuses on interaction rather than knowledge generation.
@@ 5.3.3. Study Design and Data Preparation
A within-subject design was used to compare the knowledge generation quality across different pipelines. Each participant rated knowledge generated from 9 moments, with 3 knowledge items per moment (one from each pipeline), totaling 27 knowledge items per participant.
@@ 5.3.5. Results
This underscores the need for AI systems to dynamically infer user interests and intentions from real-time context, as users may not explicitly define all their fine-grained desires in advance. It also highlights the importance of coordinating rules with user profiles.
@@ 6.1. Participants
The study consisted of two phases. In Phase 1, 18 participants (8 females, 10 males, Age: $M=23.9,\>SD=3.7$) from the university community engaged in a one-time use of AiGet. Among these participants, six reported using AI tools (e.g., Google Lens, ChatGPT app) to identify surrounding objects in their daily activities at least three times per week. Eleven participants used such tools less frequently, ranging from once per month to twice per week. The remaining participant rarely used these tools. All participants self-reported a minimum level of professional working fluency in English.
In Phase 2, 6 participants (3 females, 3 males) from Phase 1 were selected for continued usage based on their demographic diversity and availability (see Appendix D.2 for details). These participants, representing a range of self-reported engagement with their surroundings—3 with high engagement and 3 with lower engagement—took part in additional experiments involving multiple sessions.
@@ 7. Overall Discussion
In the following discussion, we will highlight two key aspects: 1) how wearable knowledge discovery assistants (i.e., wearable AI assistants supporting knowledge discovery) can be used for informal learning in daily life, and 2) what improvements are needed for their long-term usage. By examining these aspects, we aim to provide insights into the future development and integration of context-aware wearable knowledge discovery assistants in everyday contexts, further improving how we acquire and interact with information from our in-situ environment.
@@ 7.1.3. Improving User Trust for “Surprising” Content
Enhancing user trust in wearable knowledge discovery assistants is crucial, particularly when the system presents surprising or counterintuitive information. As P1 noted, prior negative experiences with AI-generated content (e.g., hallucinations (Bang et al., 2023)) can cause users to distrust unexpected information, even when it is accurate—such as learning that ‘a palm tree is more closely related to grass than typical trees’.
@@ 7.2.1. Balancing Knowledge Diversity and Depth
In the real-world study, participants highlighted AiGet’s ability to “break out of information cocoons [exposed only to information that aligns with their existing beliefs or interests] (P3)” by providing information on both prior interested topics and relatively unfamiliar content embedded in the environment, thus expanding their “knowledge base (P18)”. While this feature fosters the discovery of new interests, it also raises concerns about potential information overload. As P1 noted, “I don’t need to be an expert on everything,” emphasizing the need for careful design in content delivery, especially for repeated exposure to entities of limited interest.
@@ 7.2.3. Considering Interruptibility in Real-Time Mode and Supporting Retrospective Mode
To reduce the interference to primary tasks, beyond providing easy-to-digest content and enabling manual output adjustment, the future system could further predict user interruptibility (Pejovic and Musolesi, 2014) and automatically reduce its frequency or turn off during moments when users prefer fewer interruptions (as shown in Figure 2-Gray parts). This feature can be implemented by 1) Using MLLMs for FPV analysis and withholding notifications when users are engaged in high-concentration or time-sensitive activities (e.g., crossing a road, focusing on work, and rushing in train stations); 2) Incorporating additional input modalities, such as Electrodermal Activity signals (Tan et al., 2021), to discern variations in mental or emotional states even within the same activity, allowing nuanced adjustments (e.g., distinguishing whether a student is thinking of homework after class or enjoying a leisurely walk home); and 3) Adapting interruptibility rules based on individual preferences and interaction history.
Complementing this real-time adaptation, participants (e.g., P3) suggested incorporating a retrospective mode to catch up on interesting or useful information they might have missed while the system is in “Do Not Disturb” status. This retrospective mode could also encourage reflection on daily experiences, offering meta-cognitive insights excluded in real-time mode due to the high cognitive load involved. Additionally, it could converse with users during this period and dynamically update user preferences in the profile based on their reflections on interactions throughout the day.
@@ 8. Limitations
While our evaluation provided initial insights into using a proactive wearable knowledge discovery assistant, it had several limitations that future research should address. First, our study focused on low-demand daily activities, such as casual walking and shopping. It remains unclear whether the designed mitigations for reducing task disruption will be as effective in higher-demand scenarios, such as time-sensitive tasks. Future research should explore how proactive knowledge discovery assistants perform under varying cognitive and temporal constraints. Additionally, the study primarily involved university students and staff, who may represent early adopters of smart glasses technology. Future studies should include a more diverse demographic to ensure broader applicability. While we conducted multi-session studies over a week, longer-term evaluations are necessary to better understand the system’s impact on learning behaviors and retention. Moreover, due to hardware constraints, participants could not use the system continuously in their daily lives without being observed, potentially affecting the naturalness of their interactions. To address these gaps, future research should conduct extended, real-world deployments to capture more authentic usage patterns and assess long-term effects on learning and behavior.
@@ 9. Conclusion
We explored the feasibility of a wearable AI assistant, AiGet, that proactively analyzes users’ in-situ environments to provide personalized, context-relevant, yet often overlooked knowledge with minimal disruption to primary tasks. Through in-lab and real-world evaluations, we demonstrated AiGet’s utility in uncovering overlooked interests, enhancing task enjoyment, reviving curiosity, and deepening connections with the environment. Our findings highlight the potential of AI-assisted informal learning to transform everyday moments into enriching learning experiences. We also identified key design considerations for future systems, including balancing knowledge diversity and depth, enhancing coherence between learning moments, and supporting real-time and retrospective interaction modes. We have open-sourced this project at: https://github.com/Synteraction-Lab/AiGet and welcome contributions from the community to enhance its capability. As wearable AI assistants continue to evolve, AiGet represents a promising initial step toward seamlessly integrating informal learning into daily life, opening new avenues for ubiquitous knowledge discovery and personal growth.
Acknowledgements.
This research is supported by the National Research Foundation Singapore and DSO National Laboratories under the AI Singapore Programme (Award Number: AISG2-RP-2020-016). The CityU Start-up Grant (9610677) also provides partial support. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the author(s) and do not reflect the views of the National Research Foundation, Singapore. We extend our gratitude to the Lee Kong Chian Natural History Museum and NUS Museum for their invaluable assistance with our user studies and to all members of the Synteraction Lab for their help in completing this project. We also thank the reviewers for their valuable feedback.
@@ A.2. Knowledge Categories and Examples
This table defines four knowledge types with examples from a supermarket beverage scene. Factual Knowledge includes specific details, like product ingredients (e.g., ”Apple juice contains natural antioxidants”). Conceptual Knowledge explains relationships, such as nutritional comparisons (e.g., ”Apple juice is sweeter, orange juice has more vitamin C”). Procedural Knowledge covers methods, like creating a mocktail (e.g., ”Mix apple juice, orange juice, and ginger ale”). Metacognitive Knowledge involves self-reflection, such as understanding personal decision-making criteria.
Knowledge Type Definition Examples
Factual Knowledge Knowledge of terminology and specific details and elements “This apple juice contains natural antioxidants called polyphenols, which can help protect your cells from damage and may even support heart health”
@@ B.1. FPV Difference Calculation
B.2. User Profile Sample
A sample user profile is as follows: {”Values/Interest”: [”healthy life”, ”fitness”, ”flowers”, ”coffee lover”, ”cat”, ”fun facts/history”, ”design”], ”Age”: ”30”, ”Gender”: ”female”, ”Citizenship”: ”XX”, ”Residence”: ”XX”, ”Education”: ”PhD in visual design”, ”Occupation”: ”senior student in XX University”}.
B.3. Parallel Processing during Output Transformation